OpenStudy (anonymous):

Find all solutions in the interval [0 degrees, 360 degrees): cot^4(2x)=9

OpenStudy (anonymous):

\[\begin{align*}\cot^42x&=9\\ \cot^42x-9&=0\\ \left(\cot^22x-3\right)\left(\cot^22x+3\right)&=0\end{align*}\] So now it's a matter of solving the two equations, \[\cot^22x-3=0\\ \cot^22x+3=0\] \[\cot^22x=\pm3~~\iff~~\cot2x=\pm\sqrt3~~\text{(because the negative 3 is extraneous)}\] So for which angles does cotangent give a positive and negative \(\sqrt3\)?